Graph the system below and write it's solution.
y=1/2x-1
2x+y=-6

Answer by mananth(16946)

y=(1/2)x-1
Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 1 2 -2
y= -1 -0.5 0 -2
(x,y) ( 0 , -1 ),( 1 , -0.5 ),( 2 0 )
y=-2x-6
y= -2 x + -6

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 1 2 -2
y= -6 -8 -10 -2
(x,y) ( 0 , -6 ),( 1 , -8 ),( 2 -10 )